OpenStudy (anonymous):

integrate (sin(3x))^4

OpenStudy (experimentx):

int ((1-cos(6x))/2)^2*dx int ((1 - 2*cos(6x) + cos(6x)^2)/4) dx int 1/4 dx -int cos(6x)/2 dx + int ((1+cos(12x))/8)dx x/4 - sin(6x)/12+x/8+sin(12x)/96+k ....

OpenStudy (anonymous):

so what is the process to the answer...

OpenStudy (callisto):

@experimentX (sin(3x))^4 = [(sin3x)^2]^2 = (2cos6xsin6x)^2 ... how come it becomes ((1-cos(6x))/2)^2 ?

OpenStudy (anonymous):

can you somehow draw it because this is very confusing to read

OpenStudy (experimentx):

cos 2x = cos(x)^2 - sin(x)^2 or, cos 2x = 1- 2sin(x)^2 so sin(x)^2 = (1-cos2x)/2

OpenStudy (anonymous):

how did you get the denominators to be 12 and 96 for the fractions

OpenStudy (callisto):

 similar for 96